what are the values of x and y if t(-2,7) (x,y)= (3,-1)?
Answer by ikleyn(52756)

.

The translation  t(-2,7)  acts in the way that


   t(-2,7) (x,y) = (x-2,y+7).


Therefore, the given part means that


    x-2 = 3,

    y+7 = -1.


It implies that  x = 3 + 2 = 5;  y = -1 - 7 = -8.


ANSWER.  x = 5;  y = -8.

Solved.